KyledaGreat:

Consider the following relation. y=x−3 Find the domain and range of the inverse. Express your answer in interval notation.

KyledaGreat:

\[(-\infty,\infty)\] \[(-\infty,\infty)\]

KyledaGreat:

Are these right ?

Tranquility:

Yes

Tranquility:

That is correct